# Rethinking Neural Combinatorial Optimization for Vehicle Routing Problems with Different Constraint Tightness Degrees

### v2

@@ Abstract
Recent neural combinatorial optimization (NCO) methods have shown promising problem-solving ability without requiring domain-specific expertise. Most existing NCO methods use training and testing data with a fixed constraint value and lack research on the effect of constraint tightness on the performance of NCO methods. This paper takes the capacity-constrained vehicle routing problem (CVRP) as an example to empirically analyze the NCO performance under different tightness degrees of the capacity constraint. Our analysis reveals that existing NCO methods overfit the capacity constraint, and they can only perform satisfactorily on a small range of the constraint values but poorly on other values. To tackle this drawback of existing NCO methods, we develop an efficient training scheme that explicitly considers varying degrees of constraint tightness and proposes a multi-expert module to learn a generally adaptable solving strategy. Experimental results show that the proposed method can effectively overcome the overfitting issue, demonstrating superior performances on the CVRP and CVRP with time windows (CVRPTW) with various constraint tightness degrees.
@@ 1 Introduction
The vehicle routing problem (VRP) is an important category of combinatorial optimization problems and is encountered in many real-world applications, including transportation [1], navigation [2], and robotics [3]. Due to the NP-hard nature [4], it is prohibitively expensive to obtain exact VRP solutions. Traditional methods typically employ problem-specific heuristic algorithms, which require extensive domain-specific expertise. As an alternative, neural combinatorial optimization (NCO) methods have received widespread attention recently [5]. These methods can automatically learn problem-solving strategies from data using neural networks. Some advanced NCO methods can even obtain near-optimal VRP solutions without the need for expert knowledge [6, 7, 8, 9, 10, 11, 12].
@@
Current NCO research has been deluded by the seemingly promising results obtained on instances with certain specific tightness, lacking a comprehensive study of how constraint tightness affects model performance. To fill this gap, this paper takes CVRP as an example and performs an empirical analysis of the NCO performance under different capacity constraint tightness. We evaluate three recently proposed NCO methods (i.e., POMO, LEHD, and BQ) that are trained using CVRP instances with fixed customer demands but different capacity values, and the results are presented in Table 1. As can be seen from the table, when the CVRP has extremely tight or loose capacity constraints (C=10 and 500), the optimality gaps of the NCO methods deteriorate by at least 5.5$\times$ and 1.6$\times$, demonstrating severe performance degradation. We believe that these NCO models overfit the training instances with a specific capacity tightness, i.e., C=50, and thus exhibit limited generalization ability across instances with other capacity values. In this paper, we reveal the reasons behind the performance degradation leveraging the problem similarity between CVRP and two other VRP variants, i.e., the traveling salesman problem (TSP) and the open vehicle routing problem (OVRP). More details are provided in Section 4.1.
@@
POMO 3.66% 20.26% 5.54$\times$ 34.12% 9.32$\times$
LEHD 4.22% 36.56% 8.66$\times$ 6.82% 1.62$\times$
BQ 3.25% 18.02% 5.54$\times$ 8.16% 2.51$\times$
@@
2 Preliminaries
2.1 Problem Definition
A VRP instance $S$ can be defined on a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the node set and =$\mathcal{E}=\{(v_{i},v_{j})|v_{i},v_{j}\in\mathcal{V},v_{i}\neq v_{j}\}$ denotes the edge set. The nodes from $v_{1}$ to $v_{n}$ represent $n$ customers, and some VRPs (e.g., CVRP, OVRP) contain a depot $v_{0}$ with $\mathcal{V}=\{v_{i}\}_{i=0}^{n}$.
The solution of a VRP instance is a tour $\boldsymbol{\pi}$, which is a permutation of the nodes. Given an objective function $f(\cdot)$ of solutions, solving the VRP instance is to search the optimal solution $\boldsymbol{\pi}^{*}$ with minimal objective function value (e.g., Euclidean distance of the tour) under certain problem-specific constraints, (e.g., visiting each node exactly once in solving the traveling salesman problem (TSP)). Therefore, a VRP can be formulated as follows:
TSP
CVRP
OVRP
2.2 VRP Solution Construction
3 NCO Performance on CVRP with Different Constraint Tightnesses
3.1 Experiment Details
Dataset
Following the data generator in prior works [13, 6], we generate seven test datasets, each comprising 10,000 CVRP100 instances. Customer demands are drawn uniformly from $\{1,\dots,9\}$, so the average customer demand is 5. Vehicle capacities for these datasets are set to 10, 50, 100, 200, 300, 400, and 500, respectively. C=500 represents a loose constraint that expectedly allows a single vehicle to feasibly serve all customers in one single route, while C=10 represents the tight constraint that allows a vehicle to serve only two customers on average. All node coordinates are uniformly sampled within a unit square, so these datasets differ solely in capacity values.
NCO Methods
Metrics & Inference
@@ 3.2 Results and Observations
The experimental results are shown in Table 2. We observe that all NCO methods achieve low optimality gaps on test instances with the same constraint tightness (e.g., C=50) as the training instances. However, their performance drops significantly on instances with unseen loose or tight constraints. For example, the optimality gap of BQ model on the dataset with C=50 is 3.25%, but the gap on the dataset with C=10 drastically expands to 18.02%. POMO also shows a similar trend, with the gap on the dataset with C=50 being 3.66%, and expanding to 34.12% on the dataset with C=500. Overall, compared to the test dataset with the same capacity value as the training set, all evaluated NCO models exhibit at least a 1.7$\times$ increase in the average optimality gap (i.e., INViT has an average gap of 13.28%, which is 1.7$\times$ larger than 7.85% on the dataset with C=50) when tested on the datasets with different capacities. These results indicate that current NCO models generally overfit to the trained constraint tightness, which causes the learned strategies to be poorly applied to instances with other constraint tightnesses, ultimately significantly degrading their performance.
@@ 4 Methodology
Figure 1: Impact of Capacity on Problem Similarity. The CVRP100 solution closely resembles the OVRP solution when capacity is very small (e.g., C=10) and becomes equivalent to the TSP solution when capacity is extremely large (e.g., C=500). We systematically quantify the problem similarity in Appendix B.
@@ 4.1 Limitations of NCO Methods
When the constraint tightness changes drastically, the optimal solution structure of CVRP undergoes significant transformations, indicating that the problem requires fundamentally different solving strategies. As shown in Figure 2, under the tight capacity constraint C=10, where the capacity is capped at twice the average customer demand, each sub-tour in the CVRP solution can only serve a limited number of customers (e.g., two or three). In this case, CVRP solutions exhibit structural similarities to OVRP solutions under identical capacity constraints, as sub-tours in their solutions only visit very few customers. That is, when vehicle capacity is severely limited, the problem-solving strategy for CVRP becomes similar to that of OVRP (except that returning to the depot is neglected). Conversely, under loose capacity constraints (e.g., C=500, approaching total customer demand), the vehicles can serve all customers in a few or even a single sub-tour. The solution structure of CVRP resembles that of TSP in this situation. Therefore, TSP-oriented solving strategies could work well for solving CVRP instances with large capacities.
@@ 4.2 Training with Varying Constraint Tightness
In order to address the overfitting issue of existing NCO methods to enable the model to learn effective problem-solving strategies across different degrees of constraint tightness, we propose a straightforward yet effective training scheme, which explicitly exposes the model to the full range of constraint tightnesses. Specifically, let $C_{min}$ and $C_{max}$ denote the minimum and maximum constraint tightness degrees in the target domain (e.g., vehicle capacities in VRPs). During training, the constraint tightness degrees are uniformly sampled from $[C_{min},C_{max}]$ for each training instance. By continuously changing the constraint tightness of the training data, the NCO model is able to learn to adjust its decision-making strategy in response to varying problem domains, improving its robustness in solving general VRP instances.
In addition, we attempt to assign all instances in a training batch the identical tightness degree, randomly sampled from $[C_{min},C_{max}]$. This batch-level tightness assignment, compared to the above instance-level assignment, leads to inferior model performance. A More detailed discussion with experiments can be found in Appendix D.
@@ 4.3 Multi-Expert Module
Given the outstanding performance of the Lightweight Encoder-Heavy Decoder (LEHD) model [14, 15], we adopt it as the backbone model to exemplify the implementation of the multi-expert module. In general, LEHD uses $L$ stacked attention layers in the decoder to construct a VRP solution step by step. As shown in Figure 3, we integrate the multi-expert module after the stacked attention layers in the decoder of LEHD. The multi-expert module consists of an instance-level gate mechanism $G(\cdot)$ and $m$ expert layers $\{E_{1},E_{2},\cdots,E_{m}\}$.
@@
Given an instance, the gate mechanism assigns the node embedding matrix $H_{t}$ to the appropriate expert layer based on the constraint tightness $C_{k}$. Specifically, the $i$-th expert layer ($i\in\{1,2,\cdots,m\}$) processes instances with constraint tightnesses in the interval $[(i-1)\beta,i\beta]$, where $\beta=\frac{|C_{max}-C_{min}|}{m}$. Therefore, the gate mechanism can be formulated as a one-hot vector
@@
Each expert layer transforms the node embedding matrix $H_{t}$ into the probability matrix for node selection. To enhance the adaptability to varying constraint tightness, each expert layer comprises $m_{e}$ stacked attention layers (see Appendix A.1 for details) before the probability prediction. The calculation process within the $i$-th expert layer can be formulated as:
@@
Table 3: Comparison results between our model and the existing best model in each dataset. Gaps on C=10, C=50/100, and C=200 to 500 come from MDAM, BQ, and LEHD respectively.
@@ 5 Experiment
For comparative analysis, we evaluate the metric of optimality gap (Gap) across all methods and datasets, as done in Section 3.1. All models use the greedy search for inference. All experiments, including training and testing, are executed on a single NVIDIA GeForce RTX 3090 GPU with 24GB memory.
@@ 5.1 Comparison Results
The experimental results are presented in Table 3, which demonstrate that our method outperforms state-of-the-art NCO methods overall. Specifically, our method achieves an average gap of 1.86%, significantly lower than existing NCO models’s average best gap of 3.71%. While our method exhibits marginally inferior performance compared to the baselines at C=50 and 100, it outperforms all competing methods in other scenarios (C=10, C=200, and above). Notably, for instances with C=300 to C=500, our method maintains an optimality gap below 1%. These results highlight the effectiveness of our method in solving problems across varying degrees of constraint tightness.
@@ 5.2 Ablation Study
We conduct an ablation study to evaluate the contributions of two key components: 1) varying constraint tightness training (VCT) and 2) the multi-expert module (MEM). We compare three training configurations: 1) the model trained on CVRP100 instances with fixed capacity (C=50) and without MEM; 2) the model trained on instances with capacities $C\sim\textbf{Unif}([10,500])$ and without MEM; 3) the model trained on instances with capacities $C\sim\textbf{Unif}([10,500])$ and with MEM. Results of these models on the test sets are presented in Table 4.
From these results, we can observe that varying constraint tightness training leads to significant improvements in the model’s problem-solving ability on 5 out of 7 datasets, reducing the average gap from 10.78% to 2.41%. However, this approach causes a slight performance degradation on the datasets with C=50 and C=100. The reason may be that the different constraint tightness degrees create conflicting optimization directions, leading to suboptimal learning outcomes. When the multi-expert module is incorporated, the model achieves consistent performance gains across all datasets, reducing the average gap by 23%. These results demonstrate that while varying constraint tightness training enhances cross-constraint-tightness generalization, the integration of the multi-expert module further strengthens this capability, enabling more robust adaptation to problem instances with varying constraint tightness.
@@ 6 Versatility
To demonstrate that our proposed method is applicable to other routing problems, we validate it on CVRPTW instances with varying degrees of time window constraint tightness. We provide the CVRPTW definition and introduce a coefficient $\alpha$ to quantify the tightness of the time window constraint coefficient in Appendix E.
We apply our proposed method to the classic POMO model to solve CVRPTW instances with varying degrees of time window constraint tightness. The hyperparameter settings of the POMO model follow the description in its original literature [6]. We train the model for 1,000 epochs using reinforcement learning, with each epoch containing 10,000 samples and a batch size of 64. For training under different constraint tightness degrees, the tightness coefficient for the time window constraints is set to $\alpha\sim\text{Uniform}(0.0,3.0)$, where $\alpha\to 0.0$, $\alpha=1.0$, and $\alpha=3.0$ denotes the extremely tight, moderate, and loose constraints, respectively. For the multi-expert module incorporated into POMO, the $i$-th expert layer is specialized in solving instances with time window tightness in the range $[(i-1),i]$. Implementation details for the multi-expert module can be found in Appendix F. We train three POMO models separately: 1) the original POMO model trained with the instances of default time window constraint (i.e., $\alpha=1.0$), 2) the model with varying constraint tightness training, and 3) the model with varying constraint tightness training and the multi-expert module. These models use greedy search for inference and are evaluated on datasets with tightness coefficients ranging from $\alpha=0.2$ to $\alpha=3.0$. The experimental results are presented in Table 5.
From the experimental results, we observe that when $alpha$ changes from $1.0$ to 0.2/3.0, the optimality gap of the original POMO model increases from 11.39% to 14.95% and 33.56%, respectively. The introduction of varying constraint tightness training improves the model’s performance in all cases except for $\alpha=1.0$. The augment of the multi-expert module further enhances the model performance across all degrees of constraint tightness. These results indicate that in CVRPTW, the NCO method also has the overfitting issue and exhibits significant performance degradation when the constraint tightness changes drastically, and the proposed method can effectively tackle the limitation, enhancing NCO model for VRP with varying domains of time windows.

### v3

@@ Abstract
Recent neural combinatorial optimization (NCO) methods have shown promising problem-solving ability without requiring domain-specific expertise. Most existing NCO methods use training and testing data with a fixed constraint value and lack research on the effect of constraint tightness on the performance of NCO methods. This paper takes the capacity-constrained vehicle routing problem (CVRP) as an example to empirically analyze the NCO performance under different tightness degrees of the capacity constraint. Our analysis reveals that existing NCO methods overfit the capacity constraint, and they can only perform satisfactorily on a small range of the constraint values but poorly on other values. To tackle this drawback of existing NCO methods, we develop an efficient training scheme that explicitly considers varying degrees of constraint tightness and propose a multi-expert module to learn a generally adaptable solving strategy. Experimental results show that the proposed method can effectively overcome the overfitting issue, demonstrating superior performance on the CVRP and CVRP with time windows (CVRPTW) with various constraint tightness degrees. The code is available at https://github.com/CIAM-Group/Rethinking_Constraint_Tightness.
@@ 1 Introduction
The vehicle routing problem (VRP) is an important category of combinatorial optimization problems and is encountered in many real-world applications, including transportation [1], navigation [2], and robotics [3]. Due to the NP-hard nature [4], it is prohibitively expensive to obtain exact VRP solutions. Traditional methods typically employ problem-specific heuristic algorithms, which require extensive domain-specific expertise. As an alternative, neural combinatorial optimization (NCO) methods have received widespread attention recently [5, 6, 7, 8]. These methods can automatically learn problem-solving strategies from data using neural networks. Some advanced NCO methods can even obtain near-optimal VRP solutions without the need for expert knowledge [9, 10, 11, 12, 13, 14, 15, 16, 17].
@@
[…]
@@
[…]
@@
2 Preliminaries
2.1 Problem Definition
A VRP instance $S$ can be defined on a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the node set and $\mathcal{E}=\{(v_{i},v_{j})|v_{i},v_{j}\in\mathcal{V},v_{i}\neq v_{j}\}$ denotes the edge set. The nodes from $v_{1}$ to $v_{n}$ represent $n$ customers, and some VRPs (e.g., CVRP, OVRP) contain a depot $v_{0}$ with $\mathcal{V}=\{v_{i}\}_{i=0}^{n}$. In this work, we assume an unlimited fleet of homogeneous vehicles starting from and returning to the depot to serve all customers. This follows the classical VRP formulation used in seminal NCO papers [18, 9].
The solution of a VRP instance is a tour $\boldsymbol{\pi}$, which is a permutation of the nodes. Given an objective function $f(\cdot)$ of solutions, solving the VRP instance is to search for the optimal solution $\boldsymbol{\pi}^{*}$ with minimal objective function value (e.g., Euclidean distance of the tour) under certain problem-specific constraints (e.g., visiting each node exactly once in solving the traveling salesman problem (TSP)). Therefore, a VRP can be formulated as follows:
TSP
CVRP
OVRP
2.2 VRP Solution Construction
2.3 Definition of Constraint Tightness
For CVRP, the “constraint tightness” is directly represented by the vehicle capacity value, $C$. A tight constraint corresponds to a low capacity value (e.g., $C=10$ in our paper). In this scenario, each vehicle can only serve a very limited number of customers, forcing the solution to be composed of many short sub-tours. A loose constraint corresponds to a high capacity value (e.g., $C=500$). Here, the capacity is so large that it is barely a limiting factor, allowing a single vehicle to serve many or all customers. This makes the problem characteristics approach those of the Traveling Salesman Problem (TSP). Therefore, in the context of CVRP, constraint tightness is inversely proportional to the vehicle capacity $C$.
We also provide a more general definition of constraint tightness independent of the problem features, such as capacity, with details presented in Appendix J
3 NCO Performance on CVRP with Different Constraint Tightnesses
3.1 Experiment Details
Dataset
Following the data generator in prior works [18, 9], we generate seven test datasets, each comprising 10,000 CVRP100 instances. Customer demands are drawn uniformly from $\{1,\dots,9\}$, so the average customer demand is 5. Vehicle capacities for these datasets are set to 10, 50, 100, 200, 300, 400, and 500, respectively. C=500 represents a loose constraint that is expected to allow a single vehicle to feasibly serve all customers in a single route, while C=10 represents the tight constraint that allows a vehicle to serve only two customers on average. All node coordinates are uniformly sampled within a unit square, so these datasets differ solely in capacity values.
NCO Methods
Metrics & Inference
@@ 3.2 Results and Observations
[…]
Figure 1: Impact of Capacity on Problem Similarity. The CVRP100 solution closely resembles the OVRP solution when capacity is very small (e.g., C=10) and becomes equivalent to the TSP solution when capacity is extremely large (e.g., C=500). We systematically quantify the problem similarity in Appendix C.
@@ 4.1 Limitations of NCO Methods
When the constraint tightness changes drastically, the optimal solution structure of CVRP undergoes significant transformations, indicating that the problem requires fundamentally different solving strategies. As shown in Fig. 2, under the tight capacity constraint C=10, where the capacity is capped at twice the average customer demand, each sub-tour in the CVRP solution can only serve a limited number of customers (e.g., two or three). In this case, CVRP solutions exhibit structural similarities to OVRP solutions under identical capacity constraints, as sub-tours in their solutions only visit very few customers. That is, when vehicle capacity is severely limited, the problem-solving strategy for CVRP becomes similar to that of OVRP (except that returning to the depot is neglected). Conversely, under loose capacity constraints (e.g., C=500, approaching total customer demand), the vehicles can serve all customers in a few or even a single sub-tour. The solution structure of CVRP resembles that of TSP in this situation. Therefore, TSP-oriented solving strategies could work well for solving CVRP instances with large capacities.
@@ 4.2 Training with Varying Constraint Tightness
In order to address the overfitting issue of existing NCO methods to enable the model to learn effective problem-solving strategies across different degrees of constraint tightness, we propose a straightforward yet effective training scheme, which explicitly exposes the model to the full range of constraint tightnesses. Specifically, let $C_{min}$ and $C_{max}$ denote the minimum and maximum constraint tightness degrees in the target domain (e.g., vehicle capacities in VRPs). During training, the constraint tightness degrees are uniformly sampled from $C_{min},C_{max}]$ for each training instance. By continuously changing the constraint tightness of the training data, the NCO model is able to learn to adjust its decision-making strategy in response to varying problem domains, improving its robustness in solving general VRP instances.
In addition, we attempt to assign all instances in a training batch the identical tightness degree, randomly sampled from $C_{min},C_{max}]$. This batch-level tightness assignment, compared to the above instance-level assignment, leads to inferior model performance. A more detailed discussion with experiments can be found in Appendix E.
@@ 4.3 Multi-Expert Module
Given the outstanding performance of the Lightweight Encoder-Heavy Decoder (LEHD) model [19, 20], we adopt it as the backbone model to exemplify the implementation of the multi-expert module. In general, LEHD uses $L$ stacked attention layers in the decoder to construct a VRP solution step by step. As shown in Fig. 3, we integrate the multi-expert module after the stacked attention layers in the decoder of LEHD. The multi-expert module consists of an instance-level gate mechanism $G(\cdot)$ and $m$ expert layers $\{E_{1},E_{2},\cdots,E_{m}\}$.
@@
Given an instance, the gate mechanism assigns the node embedding matrix $H_{t}$ to the appropriate expert layer based on the constraint tightness $C_{k}$. Specifically, the $i$-th expert layer ($i\in\{1,2,\cdots,m\}$) processes instances with constraint tightnesses in the interval $(i-1)\beta,i\beta]$, where $\beta=\frac{|C_{max}-C_{min}|}{m}$. Therefore, the gate mechanism can be formulated as a one-hot vector
@@
Each expert layer transforms the node embedding matrix $H_{t}$ into the probability matrix for node selection. To enhance the adaptability to varying constraint tightness, each expert layer comprises $m_{e}$ stacked attention layers (see Appendix B.1 for details) before the probability prediction. The calculation process within the $i$-th expert layer can be formulated as:
@@
Table 3: Comparison results between our model and the existing best model in each dataset. Gaps on C=10, C=50/100, and C=200 to 500 come from MDAM, BQ, and LEHD, respectively.
@@ 5 Experiment
For comparative analysis, we evaluate the metric of optimality gap (Gap) across all methods and datasets, as done in Section 3.1. All models use the greedy search for inference. The inference time that quantifies a method’s computational efficiency is shown in Appendix A. All experiments, including training and testing, are executed on a single NVIDIA GeForce RTX 3090 GPU with 24GB of memory.
@@ 5.1 Comparison Results
The experimental results are presented in Table 3, which demonstrate that our method outperforms state-of-the-art NCO methods overall. Specifically, our method achieves an average gap of 1.86%, significantly lower than existing NCO models’ average best gap of 3.71%. While our method exhibits marginally inferior performance compared to the baselines at C=50 and 100, it outperforms all competing methods in other scenarios (C=10, C=200, and above). Notably, for instances with C=300 to C=500, our method maintains an optimality gap below 1%. These results highlight the effectiveness of our method in solving problems across varying degrees of constraint tightness.
@@ 5.2 Ablation Study
We conduct an ablation study to evaluate the contributions of two key components: 1) varying constraint tightness training (VCT) and 2) the multi-expert module (MEM). We compare three training configurations: the model trained on CVRP100 instances with fixed capacity (C=50) 1) without MEM and 2) with MEM; 3) the model trained on instances with capacities $C\sim\textbf{Unif}([10,500])$ and without MEM; 4) the model trained on instances with capacities $C\sim\textbf{Unif}([10,500])$ and with MEM. We maintain a consistent training data budget across all experimental settings. Results of these models on the test sets are presented in Table 4.
From these results, we can observe that while adding only the MEM reduces the average gap from 10.78% to 8.95%, providing a modest benefit, varying constraint tightness training leads to significant improvements in the model’s problem-solving ability on 5 out of 7 datasets, reducing the average gap from 10.78% to 2.41%. However, this approach causes a slight performance degradation on the datasets with C=50 and C=100. The reason may be that the different constraint tightness degrees create conflicting optimization directions, leading to suboptimal learning outcomes. When the multi-expert module is incorporated, the model achieves consistent performance gains across all datasets, reducing the average gap by 23% (from 2.41% to 1.86%). These results demonstrate that while varying constraint tightness training enhances cross-constraint-tightness generalization, the integration of the multi-expert module further strengthens this capability, enabling more robust adaptation to problem instances with varying constraint tightness.
@@ 6 Versatility
To demonstrate that our proposed method is applicable to other routing problems, we validate it on CVRPTW instances with varying degrees of time window constraint tightness. We provide the CVRPTW definition and introduce a coefficient $\alpha$ to quantify the tightness of the time window constraint coefficient in Appendix F.
We apply our proposed method to the classic POMO model to solve CVRPTW instances with varying degrees of time window constraint tightness. The hyperparameter settings of the POMO model follow the description in its original literature [9]. We train the model for 1,000 epochs using reinforcement learning, with each epoch containing 10,000 samples and a batch size of 64. For training under different constraint tightness degrees, the tightness coefficient for the time window constraints is set to $\alpha\sim\text{Uniform}(0.0,3.0)$, where $\alpha\to 0.0$, $\alpha=1.0$, and $\alpha=3.0$ denote the extremely tight, moderate, and loose constraints, respectively. For the multi-expert module incorporated into POMO, the $i$-th expert layer is specialized in solving instances with time window tightness in the range [(i1),i][(i-1),i]. Implementation details for the multi-expert module can be found in Appendix G. We train three POMO models separately: 1) the original POMO model trained with the instances of default time window constraint (i.e., $\alpha=1.0$), 2) the model with varying constraint tightness training, and 3) the model with varying constraint tightness training and the multi-expert module. These models use greedy search for inference and are evaluated on datasets with tightness coefficients ranging from $\alpha=0.2$ to $\alpha=3.0$. The experimental results are presented in Table 5.
From the experimental results, we observe that when alphaalpha changes from $1.0$ to 0.2/3.0, the optimality gap of the original POMO model increases from 11.39% to 14.95% and 33.56%, respectively. The introduction of varying constraint tightness training improves the model’s performance in all cases except for $\alpha=1.0$. The augmentation of the multi-expert module further enhances the model performance across all degrees of constraint tightness. These results indicate that in CVRPTW, the NCO method also has the overfitting issue and exhibits significant performance degradation when the constraint tightness changes drastically, and the proposed method can effectively tackle the limitation, enhancing the NCO model for VRP with varying domains of time windows.